# Raman Microspectroscopy for Real-Time Structure Indicator in Ultrafast Laser Writing

### v1

@@ 2 Raman Monitoring of Electrical Performance
Ultimately, the optimal feedback signal for in-situ monitoring of fabrication should be directly tied to the functional performance of the resulting devices. For graphite electrodes on diamond, this performance is best represented by the conductivity of the laser-written features. To this end, the electrical conductivity of the pads and wires was measured using a custom setup (see Electrical conductivity measurement).
@@
Figure 4 presents our findings and consolidates the link between laser scan speed, local bonding state, and device-level conductivity of the laser-written features. Figure 4a depicts the current–voltage traces for all wires and pads, which are linear over mV, indicating ohmic behavior and no measurable Schottky barriers; conduction is therefore governed by the volume fraction and connectivity of the laser-generated graphitic phase. For graphitic structure of width of 200 m written at ms-1, the pad resistance is , versus for the wire at the same speed. The lower pad resistance reflects its larger cross-section and additional over-writing during fabrication, which further graphitizes the diamond and increases the graphite fraction. Figure 4b generalises this comparison across scan speeds: the resistance decreases as the scan speed is reduced, with pad values consistently below wire values at identical speeds. All resistances are extracted from linear fits to quasi-static I–V sweeps within mV (details of the IV curves are shown in Supplementary Note 7).
@@
In Figure 4c and d we plot, separately for wires and pads, the mean normalized integrated sp3 and sp2 intensities as well as their ratio versus laser scan speed. As the scan speed increases, the resistance rises concurrently with the normalized sp3 intensity, indicating reduced graphitization and weakened conductive pathways. For wires (Figure 4c), the normalized sp2 intensity increases significantly as the scan speed decreases from high to intermediate values, corresponding to a sharp drop in resistance. However, once the sp3 contribution is strongly diminished (deeply graphitized regime), sp2 intensity alone becomes a less reliable measure of further conductivity enhancement. Since the sp2 intensity rises as the sp3 intensity decreases with scan speed, their ratio mirrors these trends. Pads show the same qualitative behavior (Figure 4d) but remain systematically more conductive than wires at identical scan speeds, consistent with their larger effective cross-section and enhanced sp2 fraction. Collectively, these data establish Raman-derived sp2/sp3 metrics as a non-contact, quantitative proxy for electrical performance and a practical handle for in-situ process control during fs-laser fabrication.
3 Discussion
We have evaluated the performance of three characterization methods (optical micrography, PL mapping, and Raman microspectroscopy) for phase transformation from insulating sp3-bonded lattice to conductive sp2-bonded carbon, an ultrafast phase change that has been extensively studied in both surface and bulk geometries [26, 4, 27]. Among these characterization methods, only Raman provides a direct, quantitative, phase-specific measurement of the bonding state.
@@ 3 Discussion
As shown in Figure 2, the sp3 and sp2 Raman features evolve with fabrication conditions and degree of graphitization. Reducing the scan speed, and thus increasing the energy deposited per unit length, leads to a steady decrease in sp3 intensity and a general increase in sp2 intensity, consistent with progressive phase conversion from diamond to graphitic carbon. At the slowest speeds, however, the sp2 intensity reaches a maximum and then declines despite continued sp3 depletion, a behavior also reported in other ultrafast graphitization studies[27, 28].
@@
This counter-intuitive reduction in sp2 intensity can be interpreted within the ‘amorphization trajectory’ framework of Ferrari et al. [29, 30], wherein progressive disorder drives a transition from nanocrystalline graphite to amorphous sp2-rich carbon and, ultimately, to highly disordered or tetrahedral amorphous forms. In this regime, shrinking and fragmenting sp2 clusters suppress well-defined Raman modes, so the apparent sp2 signal diminishes even as the actual sp2 bonding fraction remains high. At intermediate doses, ordered sp2 clusters grow, strengthening the D and G bands. At extreme doses, however, high defect densities and reduced cluster sizes suppress well-defined vibrational modes, broadening and merging the D and G peaks into a low-contrast background. The resulting spectral degradation reduces the apparent sp2 Raman intensity, even when the actual sp2 bonding fraction remains high. Under such high-dose conditions, sp3 depletion provides a more reliable metric for tracking graphitization progression, as it decreases monotonically with increasing laser dose and is insensitive to spectral contrast loss[27, 25].
@@
The stronger sp3 depletion observed for pads can be interpreted by stress-assisted graphitization during overwriting. Lattice mismatch between diamond and graphite generates tensile stress in the transformed surface layer and, crucially, in-plane compressive stress in the adjacent diamond [25]. Since the graphitization barrier decreases under lattice contraction, compression promotes earlier conversion as the lattice temperature rises. In pad writing, each new pass overlaps previously graphitized regions, introducing a local compressive field in neighboring diamond and enhancing the conversion rate. This local compressive strain field during overwriting explains the greater sp3 depletion and lower resistance of pads relative to single-pass wires at identical scan speeds, with the Raman shift direction matching established stress coefficients for diamond (blue shift in wavenumber) [31, 32, 33] (see Supplementary Note 6).
As presented in Figure 3, hyperspectral unmixing could also serve as a label-free, instrument-agnostic tool for generation of phase-fraction maps from 2D Raman data. The unmixing technique agrees with simpler ratio metrics while exploiting the full spectrum. By discovering two data-driven endmembers (ascribed to diamond and graphite phases) and enforcing non-negativity/sum-to-one constraints, the linear mixing model yields abundances that are directly interpretable as local composition and are robust to power drift, focus variations, and baseline fluorescence, issues that can affect single-band readouts. The approach is not without limits: severe disorder, stress-induced line shifts, or strong fluorescence can deviate from strict linearity and slightly bias abundance estimates.
Electrical measurements shown in Figure 4 reveal ohmic behavior for both pads and wires, with pads exhibiting systematically lower resistance due to their larger cross-section and cumulative overwriting during fabrication. Notably, the correlation between Raman metrics and resistance highlights a counter-intuitive result: sp3 depletion is a more robust indicator of electrical performance than the absolute sp2 signal. This distinction likely arises because sp3 suppression directly reflects the fraction of diamond lattice converted, irrespective of the ordering of the resulting carbon, whereas sp2 intensity is sensitive to disorder and may be suppressed even in highly conductive, amorphous-rich networks. Similar conclusions have been drawn in buried-wire studies where the attenuation of the diamond peak correlated linearly with conductive cross-section. Moreover, as shown in Figure 1, the superior SNR of the sp3 line relative to the sp2 band makes sp2-based metrics inherently more susceptible to broadband background from non-graphitic damage (see Supplementary Note 2). This is also true for full-spectrum based metrics, such as hyperspectral unmixing (see Supplementary Note 8).
From a metrology standpoint, this insight opens a streamlined path to in-situ control. In our experiments, one-second integrations of the diamond peak were sufficient to resolve sp3 depletion without sample cleaning. This can be further aided by the incidental removal of surface debris by the 532 nm probe laser (see Supplementary Note 1) which also reduces the masking effect of debris that might affect the fabrication.
III Conclusion and Outlook
Femtosecond laser writing in diamond produces conductive carbon features whose performance can be anticipated from phase-specific Raman spectral features. We have provided two-dimensional Raman mapping of laser-written surface structures under different fabrication conditions. Overall, our findings show that the sp2/sp3 intensity ratio provides a robust and reliable indicator of graphitization progression. Because it is derived from relative band intensities, this ratio is largely insensitive to variations in laser power and other experimental fluctuations. However, under high-dose irradiation conditions, monitoring sp3 depletion becomes a more consistent metric, as it decreases monotonically with increasing laser dose and remains unaffected by the loss of spectral contrast that accompanies extreme structural disorder [27, 25].
To make the readout robust, we complemented simple intensity ratios with blind hyperspectral unmixing performed without any pre-labelling. The algorithm discovers two endmembers (sp3-like and sp2-like) directly from the data and returns per-pixel abundances that agree with PL/Raman contrast while exploiting the full spectrum. This analysis also shows that comparing the background sp3 (pristine diamond) to the written sp2 provides the highest-SNR indicator of conversion and is therefore ideal for in-situ characterization. This means that a full spectrometer may not be necessary for feedback: the 1332 cm-1 line could be monitored using narrowband filtering and a CCD or APD, reducing system complexity and acquisition time. By establishing a depletion threshold corresponding to the target resistance, the fabrication process could be terminated or parameters adjusted in real time. Closed-loop optimization approaches, potentially incorporating machine learning for rapid spectral classification [34], could then maintain consistent electrical performance across large-area, multi-feature devices.
Beyond diamond and electrical performance, the same characteisation framework can be optimized and implemented in other materials by (a) acquiring (offline) representative sample datasets under the target fabrication conditions and apply the same label-free post-processing (ratio maps and blind unmixing) to discover robust, instrument-agnostic indicators and setpoints (e.g., depletion of a host phonon, bounds on D/G, endmember abundances); and (b) deploy those indicators as real-time monitors during fabrication using narrowband detection or lightweight spectral sampling. This strategy generalises as a local phase/composition sentinel across laser fabrication processes: (i) waveguide manufacturing in glasses and wide-bandgap crystals, where Raman-visible densification/strain fingerprints correlate with refractive-index change and loss [35, 36]; (ii) laser-induced graphene on polymers/resists [37], where D/G evolution guides sheet resistance and uniformity; (iii) embedded heaters/electrodes in SiC, sapphire, silica, and ceramics, where host phonons and carbon signatures define safe process windows [38, 39]; and (iv) phase-specific transformations in oxides and bio-relevant substrates, where a Raman marker provides a chemically selective stop condition [40].
In summary, hyperspectral Raman provides a phase‐specific, quantitative readout that links fabrication conditions to device performance during ultrafast laser writing in diamond. Across pads and wires, we find that depletion of the diamond sp3 line at 1332 cm-1 is a more reliable predictor of conductivity than absolute sp2 intensity, which can diminish at high disorder. A simple, self‐referenced sp3 metric, benchmarked to the local pristine background, therefore offers a robust in situ indicator, while a ratio-based metric, such as sp2-to-sp3 ratio offers better robustness to power, focus, and collection variations. Combined with label‐free unmixing to extract per‐pixel phase abundances in case of post-fabrication mapping, this framework enables calibrated thresholds for stop/adjust decisions and is readily implemented with narrowband detection for real‐time process control. Because the indicators are phase‐agnostic and instrument‐agnostic, the approach generalizes to other hosts and functionalities (e.g., waveguides, embedded heaters/electrodes, laser‐induced graphene), providing a practical route toward specification‐driven fs‐laser microfabrication.
IV Materials and Methods
1 Electrode fabrication
@@
A single-crystal, optical-grade diamond grown by chemical-vapor deposition (Element Six; nominal N concentration ppm) served as the substrate for this study. Surface graphitic pads and wires were inscribed with a fs laser operating at 520 nm ( fs, kHz) with various laser scan speed. Wave-front aberrations were corrected by a liquid-crystal spatial light modulator, and the beam was focused through an objective lens with numerical aperture (ZEISS Plan-Neofluar, 0.5NA, 20x). Wire electrodes were written in a single pass under continuous stage translation (no overwriting). Successive pulses partially overlapped to form a continuous track; at fixed the pulse-to-pulse separation was set by the scan speed. Pad electrodes were produced by stacking parallel wires with a constant 0.5 m pitch, maintained for all scan speeds.
2 Electrical conductivity measurement
@@ 3 Brightfield imaging
4 Confocal Imaging
Our fabrication setup has an integrated custom confocal microscope that can be simultaneously utilized for photoluminescence and Raman spectral collection. Using a 532 nm continuous-wave excitation laser (Cobolt Samba 150mW) and a fast steering mirror (FSM-300, Newport) relayed via a 4f into the back aperture of the system’s objective, 2D PL and RS maps can be collected. Notably, in this study, the 2D maps were collected post-fabrication utilising a higher numerical aperture objective (Olympus PlanApo 0.95NA, 60x).
Both photoluminescence and Raman spectra were collected simulataneously employing 30 mW of laser excitation. The samples were scanned with a 500nm pixel-to-pixel lateral displacement and a 1 second exposure time. The collected signal was split 99:1 between the Raman and the photoluminescence collection paths. For the latter, signal was collected via a single photon avalanche detector (SPCM-AQRH-14-FC, Excelitas Inc.). Given that the collection path of our confocal setup features a 550 nm long-pass dichroic, and that the SPAD quantum efficiency drops significantly after 850 nm, the PL collection window can be considered 550-850 nm. For the Raman collection path, scattered light was dispersed by a 1200 lines mm-1 blazed spectrograph (SpectraPro HRS-1200) and detected on a deep-cooled CCD camera (PIXIS 100, Princeton Instruments).
5 Raman spectra preprocessing
@@
From , we computed integrated intensities for sp3 (diamond, cm-1) and sp2 (graphitic, cm-1) at each pixel and identified darkest sp3 pixels ( for wire and for pad) as modified regions. For these pixels, we report the spatial coordinates, the normalized integrals and , their averages, and the ratio as a quantitative metric of graphitization. Fixed integration windows were used throughout (sp3 width cm-1; sp2 width cm-1); thus, the sp2/sp3 ratio is the ratio of integrated band intensities, not intensity per unit wavenumber. Full algorithmic details, parameter choices, and examples are provided in the SI (see Supplementary Note 3).
6 Spectral Unmixing
@@
To identify the endmember spectra from the data, we employed two well-established blind unmixing algorithms (a)Vertex Component Analysis (VCA) [41], a fast geometrical method that projects the dataset along random directions to iteratively identify extreme points of the spectral simplex, which serve as candidate endmembers; and (2) N-FINDR [42], an algorithm that searches for the set of endmembers maximising the volume of the simplex formed by candidate spectra, under the assumption that pure pixels are present in the dataset. Both algorithms require specifying the number of endmembers in advance. Here, we constrained the problem to two endmembers (sp2 and sp3), which reflects the known physics of the system and avoids overfitting.
@@
Once endmembers are identified, abundances can computed by solving the linear inverse problem under different constraints: (a) Fully Constrained Least Squares (FCLS) [43]: enforces both non-negativity (), and the sum-to-one constraint (), ensuring abundances can be interpreted as fractions; and (b) Non-Negative Least Squares (NNLS) [44]: relaxes the sum-to-one condition but maintains non-negativity, which can be useful when baseline variations are present. As shown in Figure 3, the abundance maps at 50 m/s, the pristine diamond region does not show an abundance of exactly 0 for the graphite endmember and 1 for the diamond endmember. This is expected since the endmembers are extracted across all wire maps, focus and intensity differences between datasets also shift the estimated simplex (see Supplementary Note 8). The result is a small residual assignment of sp2 abundance even in pristine regions. Importantly, the maps still provide clear contrast: pristine areas are dominated by the sp3 endmember, while graphitized tracks are dominated by sp2, confirming the physical validity of the decomposition.
@@
Raman scattering is, to first order, linear in the number of scatterers, so spatial coexistence of sp3 and sp2 domains within the confocal volume leads to additive spectra, which the LMM captures. Moreover, our maps are recorded on polished surfaces with modest absorption at 532 nm outside the most heavily graphitized regions, reducing multiple scattering and re-absorption that could induce non-linear mixing. That said, three effects can challenge strict linearity: (i) strong broadband fluorescence in highly damaged zones, (ii) self-absorption and re-emission in thick, strongly absorbing tracks, and (iii) spectral distortions from stress/strain gradients that shift or broaden peaks in a way not representable by a fixed endmember. In such cases, non-linear or bilinear models (e.g., polynomial post–nonlinear mixing, kernel-based unmixing, or autoencoder-based unmixing [45]) may yield incremental gains, at the cost of additional parameters and reduced interpretability.
@@
We conducted the hyperspectral unmixing analysis with Python using the RamanSPy package [46].
V Acknowledgments

### v2

@@ 2.2 Raman Monitoring of Electrical Performance
Ultimately, the optimal feedback signal for in-situ monitoring of fabrication should be directly tied to the functional performance of the resulting devices. For graphite electrodes on diamond, this performance is best represented by the conductivity of the laser-written features. To this end, the electrical conductivity of the pads and wires was measured using a custom setup (see Materials and Methods).
@@
Figure 4 presents our findings and consolidates the link between laser scan speed, local bonding state, and device-level conductivity of the laser-written features. Figure 4a depicts the current–voltage traces for all wires and pads, which are linear over mV, indicating ohmic behavior and no measurable Schottky barriers; conduction is therefore governed by the volume fraction and connectivity of the laser-generated graphitic phase. For graphitic structure of width of 200 m written at ms-1, the pad resistance is , versus for the wire at the same speed. The lower pad resistance reflects its larger cross-section and additional over-writing during fabrication, which further graphitizes the diamond and increases the graphite fraction. All resistances are extracted from linear fits to quasi-static I–V sweeps within mV (details of the IV curves are shown in Supporting Note 10 and 11).
@@
In Figure 4c and d we plot, separately for wires and pads, the mean normalized integrated sp3 and sp2 intensities as well as their ratio versus laser scan speed. As the scan speed increases, the resistance rises concurrently with the normalized sp3 intensity, indicating reduced graphitization and weakened conductive pathways. For wires (Fig. 4c), decreasing the scan speed from high to intermediate increases the normalized sp2 intensity and lowers the resistance. However, once the sp3 contribution is strongly diminished (deeply graphitized regime), sp2 intensity alone becomes a less reliable measure of further conductivity enhancement. Since the sp2 intensity rises as the sp3 intensity decreases with scan speed, their ratio mirrors these trends. Pads show the same qualitative behavior (Figure 4d) but remain systematically more conductive than wires at identical scan speeds.
2.3 Discussion
@@
Figure 2 summarizes how the Raman features evolve with writing dose, controlled by scan speed. Reducing the speed, which increases the energy deposited per unit length, produces a monotonic depletion of the 1332 cm-1 sp3 line and a general rise of the 1580 cm-1 sp2 band, consistent with progressive diamond-to-graphite conversion. At the slowest speeds, however, the sp2 intensity reaches a maximum and then declines despite continued sp3 depletion, a behavior also reported in other studies on diamond graphitization [Ali2022, Khomich2023].
@@
This counter-intuitive reduction in sp2 intensity can be interpreted within the ‘amorphization trajectory’ framework of Ferrari et al. [Ferrari2000, Ferrari2004], wherein progressive disorder drives a transition from nanocrystalline graphite to amorphous sp2-rich carbon and, ultimately, to highly disordered or tetrahedral amorphous forms. In this regime, shrinking and fragmenting sp2 clusters suppress well-defined Raman modes, so the apparent sp2 signal diminishes even as the actual sp2 bonding fraction remains high. At intermediate doses, ordered sp2 clusters grow, strengthening the D and G bands. At extreme doses, however, high defect densities and reduced cluster sizes suppress well-defined vibrational modes, broadening and merging the D and G peaks into a low-contrast background. The resulting spectral degradation reduces the apparent sp2 Raman intensity, even when the actual sp2 bonding fraction remains high. Therefore, under such high-dose conditions, sp3 depletion provides a more reliable metric for tracking graphitization progression, as it decreases monotonically with increasing laser dose, is insensitive to spectral contrast loss, and, thanks to its higher SNR, is less sensitive to non-specific broadband background [Ali2022, Wang2000].
@@
Pad writing shows stronger sp3 depletion than single-pass wires, which we attribute to stress-assisted graphitization during overwriting. Lattice mismatch generates tensile stress in the transformed layer and in-plane compressive stress in adjacent diamond [Wang2000]. Compression lowers the graphitization barrier, so overlapping passes enhance conversion. The observed wavenumber shift of the diamond line is consistent with established stress coefficients [Grimsditch1975, Zaitsev2001, Orwa2000] (see Supporting Note 6) and explains the lower resistance measured for pads at identical scan speeds.
In Figure 3 we introduce the use of more advanced Raman-based structure-characterization metrics. Particularly we first present multi-band metrics, mainly sp3-to-sp2 ratio. Using this ratio has the advantage of being agnostic to some of the effects that can generate non-structure-specific intensity fluctuations in the sp2 and sp3 peaks, such as excitation power. Moreover, we also introduce hyperspectral unmixing as a label-free complementary spectral analysis (Fig. 3). A linear mixing model with non-negativity and sum-to-one constraints recovers two data-driven endmembers that map onto diamond-like and graphite-like responses, producing phase-fraction maps that agree with band-based indicators while leveraging the full spectrum. The approach is advantageous as it does not rely on pre-established bands of interest and is also robust to power drift and can potentially help reduce the effect of baseline fluorescence, though the linearity assumption presents some limitations (see Materials and Methods). Yet, in the case of monitoring graphitization in diamond, where the bands of interest are well-known, hyperspectral unmixing does not provide a significant advantage.
Finally, in Figure 4 we show how Raman spectroscopy can be used to evaluate the performance of graphite electrodes. Across devices, depletion of the 1332 cm-1 sp3 line correlates with resistance more robustly than absolute sp2 intensity. This is likely due to a combination of the sensitivity of sp2 to disorder in high-dose conditions and the effect of the broadband background from nongraphitic damage, which is more severe in sp2 in the low-dose regime due to the difference in SNR with the sp3 line (Fig. 1, Supporting Note 2). As shown in Figures 4c,d, the sp2-to-sp3 ratio is a useful indicator of graphitization because relative normalization reduces sensitivity to power fluctuations. Yet, under high-dose irradiation, sp3 exhibits less variance than sp2-to-sp3 ratio, likely due to the effect of extreme disorder on the sp2 band [Ali2022, Wang2000].
3 Conclusion and Outlook
Raman microspectroscopy provides a phase-specific, quantitative readout that links to ultrafast laser writing conditions and device performance in diamond. Across surface pads and wire electrodes, depletion of the 1332 cm-1 sp3 line predicts conductivity more reliably than absolute sp2 and PL intensities. Some of these deficiencies could be solved via spectral pre-processing, which has the drawback of adding complexity, time, and additional sources of error [chen2018fast].
Therefore, a self-referenced sp3 metric, benchmarked to the local pristine background, is therefore a robust in situ indicator, while the sp2/sp3 ratio remains useful at moderate doses since it is less sensitive to changes in experimental conditions. Thus, Raman microspectroscopy can be integrated into the fabrication system to provide real-time, in-situ feedback of the device fabrication. In this integrated application, rather than generating the maps shown in this work, the excitation laser could be spatially offset with a lag with the fabrication laser and sample scattering could be collected with sub-second integration times. Additionally, as shown in Supporting Note 2, the excitation beam would contribute to removing surface debris.
As part of this work, linear hyperspectral unmixing has been introduced as a full-spectrum, label-free alternative for post-fabrication mapping and indicator validation. In the case of diamond surface graphitization, unmixing performs analogously with band-based metrics, demonstrating its potential for samples with less defined structure-specific bands.
The methodology presented in this work is not specific to diamond structure indicator and could therefore be adapted to other hosts and functionalities (e.g., fabrication in glasses and wide-bandgap crystals [fernandez2014role, Li2008, Little2010], laser-induced graphene on polymers [Wang2025], and embedded heaters or electrodes in SiC, sapphire, silica, and ceramics [Choi2016, Naseri2022, Zhang2025]), providing a practical route toward specification‐driven fs‐laser microfabrication.
4 Materials and Methods
4.1 Electrode fabrication
@@
A single-crystal, optical-grade diamond grown by chemical-vapor deposition (Element Six; nominal N concentration ppm) served as the substrate for this study. Surface graphitic pads and wires were inscribed with a fs laser operating at 520 nm ( fs, kHz,) and a pulse energy of 30 nJ, using various laser scan speed. Wave-front aberrations were corrected by a liquid-crystal spatial light modulator, and the beam was focused through an objective lens with numerical aperture (ZEISS Plan-Neofluar, 0.5NA, 20x). Wire electrodes were written in a single pass under continuous stage translation (no overwriting). Successive pulses partially overlapped to form a continuous track; at fixed the pulse-to-pulse separation was set by the scan speed. Pad electrodes were produced by stacking parallel wires with a constant 0.5 m pitch, maintained for all scan speeds.
4.2 Electrical conductivity measurement
@@ 4.3 Brightfield imaging
4.4 Confocal Imaging
Our fabrication setup has an integrated custom confocal microscope that can be simultaneously utilized for photoluminescence and Raman spectral collection. Using a 532 nm continuous-wave excitation laser (Cobolt Samba 150mW) and a fast steering mirror (FSM-300, Newport) relayed via a 4f into the back aperture of the system’s objective, 2D photoluminescence and Raman maps can be collected. Notably, in this study, the 2D maps were collected post-fabrication utilizing a higher numerical aperture objective (Olympus PlanApo 0.95NA, 60x).
Both photoluminescence and Raman spectra were collected simultaneously employing 30 mW of laser excitation. The samples were scanned with a 500nm pixel-to-pixel lateral displacement and a 1 second exposure time. The collected signal was split 99:1 between the Raman and the photoluminescence collection paths. For the latter, signal was collected via a single photon avalanche detector (SPCM-AQRH-14-FC, Excelitas Inc.). Given that the collection path of our confocal setup features a 550 nm long-pass dichroic, and that the SPAD quantum efficiency drops significantly after 850 nm, the PL collection window can be considered 550-850 nm. For the Raman collection path, scattered light was dispersed by a 1200 lines mm-1 blazed spectrograph (SpectraPro HRS-1200) and detected on a deep-cooled CCD camera (PIXIS 100, Princeton Instruments).
4.5 Raman spectra preprocessing
@@
From , we computed integrated intensities for sp3 (diamond, cm-1) and sp2 (graphitic, cm-1) at each pixel and identified darkest sp3 pixels ( for wire and for pad) as modified regions. For these pixels, we report the spatial coordinates, the normalized integrals and , their averages, and the ratio as a quantitative metric of graphitization. Fixed integration windows were used throughout (sp3 width cm-1; sp2 width cm-1); thus, the sp2/sp3 ratio is the ratio of integrated band intensities, not intensity per unit wavenumber. Full algorithmic details, parameter choices, and examples are provided in the SI (see Supporting Note 3).
4.6 Spectral Unmixing
@@
To identify the endmember spectra from the data, we employed two well-established blind unmixing algorithms (a)Vertex Component Analysis (VCA) [nascimento2005vertex], a fast geometrical method that projects the dataset along random directions to iteratively identify extreme points of the spectral simplex, which serve as candidate endmembers; and (2) N-FINDR [winter1999n], an algorithm that searches for the set of endmembers maximising the volume of the simplex formed by candidate spectra, under the assumption that pure pixels are present in the dataset. Both algorithms require specifying the number of endmembers in advance. Here, we constrained the problem to two endmembers (sp2 and sp3), which reflects the known physics of the system and avoids overfitting.
@@
Once endmembers are identified, abundances can computed by solving the linear inverse problem under different constraints: (a) Fully Constrained Least Squares (FCLS) [heinz2001fully]: enforces both non-negativity (), and the sum-to-one constraint (), ensuring abundances can be interpreted as fractions; and (b) Non-Negative Least Squares (NNLS) [lawson1995solving]: relaxes the sum-to-one condition but maintains non-negativity, which can be useful when baseline variations are present. As shown in Figure 3, the abundance maps at 50 m/s, the pristine diamond region does not show an abundance of exactly 0 for the graphite endmember and 1 for the diamond endmember. This is expected since the endmembers are extracted across all wire maps, focus and intensity differences between datasets also shift the estimated simplex (see Supporting Note 8). The result is a small residual assignment of sp2 abundance even in pristine regions. Importantly, the maps still provide clear contrast: pristine areas are dominated by the sp3 endmember, while graphitized tracks are dominated by sp2, confirming the physical validity of the decomposition.
@@
Raman scattering is, to first order, linear in the number of scatterers, so spatial coexistence of sp3 and sp2 domains within the confocal volume leads to additive spectra, which the LMM captures. Moreover, our maps are recorded on polished surfaces with modest absorption at 532 nm outside the most heavily graphitized regions, reducing multiple scattering and re-absorption that could induce non-linear mixing. That said, three effects can challenge strict linearity: (i) strong broadband fluorescence in highly damaged zones, (ii) self-absorption and re-emission in thick, strongly absorbing tracks, and (iii) spectral distortions from stress/strain gradients that shift or broaden peaks in a way not representable by a fixed endmember. In such cases, non-linear or bilinear models (e.g., polynomial post–nonlinear mixing, kernel-based unmixing, or autoencoder-based unmixing [georgiev2024hyperspectral]) may yield incremental gains, at the cost of additional parameters and reduced interpretability.
@@
We conducted the hyperspectral unmixing analysis with Python using the RamanSPy package [georgiev2024ramanspy].
5 Acknowledgments